# LEMUR: Learned Multi-Vector Retrieval

### v1

@@ 4.1 Network architecture
We use a small two-layer MLP as the network $\phi$. This allows us to keep the inference cost trivial and to train the network quickly even on a CPU. The feature encoder $\psi$ is defined as
$\psi(x)=\text{LN}(\text{GELU}(W^{\prime}x+b)),$
where $\text{LN}(\cdot)$ is layer normalization (Ba et al., 2016) and $\text{GELU}(\cdot)$ is the GELU activation function (Hendrycks and Gimpel, 2016), $W^{\prime}\in\mathbb{R}^{d^{\prime}\times d}$, and $b\in\mathbb{R}^{d^{\prime}}$. We did not observe significant performance gains from increasing the network depth beyond two layers. Based on our ablation study (See Sec. 6.2), we fix the hidden layer size to $d^{\prime}=2048$ in all other experiments. Increasing $d^{\prime}$ further improves the accuracy of the MaxSim estimates, but has only a small impact on the end-to-end performance, since it also increases the computational cost of retrieval.
In all of the experiments, we always use the same hyperparameters to train the model; see Appendix A.
@@ 4.2 Training set selection
Since only some of the datasets used in our experiments have a separate set of training queries, we use the corpus documents encoded using the query encoder to train Lemur in all of our experiments for consistency. Formally, we sample a small subset of $n^{\prime}\ll m$ corpus documents, denoted by $\mathcal{C}_{1},\dots\mathcal{C}_{n^{\prime}}$, encode them using the query encoder of the multi-vector model as $X_{i}=Q(\mathcal{C}_{i})$ for each $i\in[n^{\prime}]$, and sample $n$ token embeddings for the training set of the model $\phi$ from the set $\cup_{i=1}^{n^{\prime}}X_{i}$.
In case there is a separate training set of queries $\mathcal{X}_{1},\dots,\mathcal{X}_{n^{\prime}}$ available, the query encoder of the multi-vector model can be used to generate the multi-vector representations $X_{1},\dots,X_{n^{\prime}}$, where $X_{i}=Q(\mathcal{X}_{i})$, and the training set can be sampled from $\cup_{i=1}^{n^{\prime}}X_{i}$. Our results (see Appendix D.2) indicate that using actual queries (when available) for training yields even higher performance than using documents.
Surprisingly, even when the training set is sampled directly from the embeddings of the corpus $C_{1},\dots,C_{m}$ (encoded using the document encoder $D$), Lemur has a consistent performance and still outperforms the baseline methods (see Appendix D.1). This training method requires neither additional data nor the use of an encoder. In summary, our results suggest that Lemur is relatively robust with respect to the type of data used to train the model.
4.3 Scalable model training
To scale the model training to large corpora, we pre-train the feature encoder $\psi$ using a subset of the documents as targets, i.e., we randomly sample $m^{\prime}\ll m$ documents and define the target function as $g^{\prime}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m^{\prime}}$,
$g^{\prime}=g_{I^{\prime}},$
where $I^{\prime}\subset[m]$ denotes the indices of the $m^{\prime}$ sampled documents. Using these auxiliary targets, we train a model $\phi^{\prime}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m^{\prime}}$,
$\phi^{\prime}(x)=W^{\prime\prime}\psi(x),$
where $W^{\prime\prime}\in\mathbb{R}^{m^{\prime}\times d^{\prime}}$. We fix the weights of the feature encoder $\psi$, and learn the $j$th row of the final weight matrix $W$ of the output layer of the final model $\phi$ as
$w_{j}=\underset{\beta\in\mathbb{R}^{d^{\prime}}}{\operatorname*{arg\,min}}\,\,E\left[\|\beta^{T}\psi(x)-g_{j}(x)\|_{2}^{2}\right]$ (7)
for each $j\in[m]$. When the weights of $\psi$ are fixed, (7) is a standard linear regression task that admits a closed form OLS solution. To speed up computation, we sample a smaller training set of size $n^{\prime}$ for computing the OLS solutions. On our experimental setup (see Sec. 6.1), we are able to process over 1000 documents per second; this also means that new documents can be indexed quickly.
@@ 5.1 Multi-vector similarity search
The retrieval system of ColBERT (Khattab and Zaharia, 2020) relies on token-level pruning of documents: For each query token, the most similar document tokens are first retrieved from $\cup_{j=1}^{m}C_{j}$, i.e., from the the set of all the document tokens. The documents containing the retrieved tokens are reranked by evaluating their MaxSim similarities with the query. PLAID (Santhanam et al., 2022a), DESSERT (Engels et al., 2023), EMVB (Nardini et al., 2024), and IGP (Bian et al., 2025) improve the latency and memory efficiency of this basic scheme by adding different pruning and approximation steps.
Like Lemur, MUVERA (Jayaram et al., 2024) reduces multi-vector similarity search to single-vector similarity search by generating single-vector encodings for the query and the documents, after which any single-vector ANNS library can be used for retrieval. However, the representations of MUVERA are data-oblivious, whereas the representations of Lemur are learned in a supervised fashion by a model that is directly optimized for the task of interest.
@@ 5.2 Reducing the storage cost of multi-vector models
Compared to single-vector retrieval, multi-vector retrieval methods have both larger memory footprints and higher latencies because of the large number of embeddings that have to be stored and processed. There exist many recent works (e.g., Lee et al., 2023; Clavié et al., 2024; MacAvaney and Tonellotto, 2024; Yan et al., 2025; Veneroso et al., 2025) that address the increased storage cost by decreasing the number of embeddings per document. These methods are orthogonal to our approach and can be combined with it.
@@ 6.1 Experimental setup
Each experiment is run on a compute node with two 20 core Intel Xeon Gold 6230 (Cascade Lake) CPUs. Each CPU has support for AVX-512 instructions and we run our experiments on all 40 cores, parallelizing over queries.
@@
Figure 2: Ablation study on the effect of the hidden layer size $d^{\prime}$ on the performance of Lemur. Left: Comparison of $\text{recall100@}k^{\prime}$ for three values of $d^{\prime}$ as a function of the candidate set size $k^{\prime}$. Right: Comparison of the end-to-end-performance comparison between different values of $d^{\prime}$ with ANNS and reranking included. While larger values of $d^{\prime}$ can yield more accurate estimates, the end-to-end performance gap is narrower due to increased ANNS complexity, yielding diminishing returns.
Datasets. For text models, we use six retrieval datasets from the BEIR (Thakur et al., 2021) benchmark. Following Santhanam et al. (2022a) and Jayaram et al. (2024), we use the development set on MS MARCO and the test set on all the other datasets. For visual document retrieval models, we pool together images and queries from five english language ViDoRe V3 (Loison et al., 2026) datasets. Table 1 lists for each dataset the number of documents, the number of test queries, and the average number of embeddings per document (using ColBERTv2 for the text datasets, and ColQwen2 for ViDoRe). We use the value $k=100$ in all of the experiments of this section; for additional results with $k=10$, see Appendix B.2.
Embedding models. In our main experiments, we use the widely-used ColBERTv2 (Santhanam et al., 2022b) model. In Sec. 6.3, we measure the end-to-end performance also on the answerai-colbert-small-v1 (Clavié, 2025), GTE-ModernColBERT-v1 (Chaffin, 2025), LFM2-ColBERT-350M (Amini et al., 2025), and jina-colbert-v2 (Jha et al., 2024) text retrieval models, and the ColModernVBERT (Teiletche et al., 2025) and ColQwen2-v1.0 (Faysse et al., 2025) visual document retrieval models.
@@ 6.2 Ablation study
Figure 3: Ablation study on the effect of using ANNS in Lemur. ANNS significantly speeds up retrieval at recall levels <0.95<0.95.
@@
[…]
We then study the effect of the latent space dimension $d^{\prime}$ of the model to the end-to-end latency (see Fig. 2, right, and Appendix B). While increasing $d^{\prime}$ from 1024 to 2048 significantly improves latency, there is a point of diminishing returns when $d^{\prime}$ is increased to 4096: even though a larger latent space dimension yields higher-quality candidates (as established in Fig. 2, left), it also increases the computational cost of querying the ANNS index, which partially offsets the accuracy gains. This also explains why the performance gap to 10240-dimensional FDEs is even larger in the end-to-end setting. While using the value $d^{\prime}=4096$ leads to lower latency than $d^{\prime}=2048$ on some of the datasets (see Appendix B), the differences in performance are small. Hence, we use $d^{\prime}=2048$ in our end-to-end experiments to decrease the memory consumption.
ANNS library vs. exact inference. Finally, we study the effect of using an ANNS index instead of evaluating the exact inner products (6) to retrieve the top-$k^{\prime}$ candidate documents. On small datasets ($m\leq 50000$), evaluating the exact inner products is sufficiently fast. However, on large datasets, an ANNS index has potential to speed up retrieval significantly. We compare exact inner product computation to approximate top-$k^{\prime}$ retrieval using the Glass (Wang, 2025) ANNS library on the NQ dataset (with $m=2681468$ documents). The results (Fig. 3) show that using an ANNS library speeds up retrieval significantly at recall levels <0.95<0.95.
Table 2: Best QPS at $\geq 80\%$ recall ($k=100$).
@@
HotpotQA 425 22 37 10
NQ 950 79 107 38 16
Quora 6503 787 679 284 89
SCIDOCS 3353 391 320 285 85
ArguAna 6924 891 467 576 76
6.3 End-to-end performance
In this section, we compare the end-to-end performance of Lemur against MUVERA (Jayaram et al., 2024), DESSERT (Engels et al., 2023), IGP (Bian et al., 2025), and PLAID (Santhanam et al., 2022a). We use recall to measure the effectivity and queries per second (QPS) to measure the efficiency of the algorithms.
Lemur implementation. For Lemur, we use PyTorch for both training and inference of the model described in Sec. 4. As a single-vector similarity search library, we use Glass (Wang, 2025), an efficient implementation of HNSW (Malkov and Yashunin, 2018) with scalar quantization. We implement the MaxSim reranking using C++.
Baseline implementations. For MUVERA, we use the official C++ implementation to generate the FDEs. To ensure a fair comparison, we combine also the FDEs of MUVERA with Glass and the same reranking implementation we use for Lemur. For DESSERT and IGP, we use the official C++ implementations, and for PLAID, we use the Rust implementation in the fast-plaid library (Sourty, 2025).
Index hyperparameters. For both Lemur and the baseline methods, we fix the index hyperparameters to values recommended by authors, except for DESSERT, for which we perform a grid search over $L\in\{32,64\},C\in\{5,7\}$. Specifically, for Lemur, we use the hidden layer size $d^{\prime}=2048$; for MUVERA, as recommended by Jayaram et al. (2024), we set $R_{\text{reps}}=40$, $k_{\text{sim}}=6,d_{\text{proj}}=d$ and then apply a final projection to generate 10240-dimensional FDEs; for PLAID, DESSERT, and IGP, we set the number of clusters as $16\sqrt{n}$ (rounded down to the nearest power of two), where $n$ is the total number of corpus token embeddings.
Query hyperparameters. For Lemur and MUVERA, we vary the candidate list size parameter (ef_search) of the ANNS library, and the number of documents reranked ($k^{\prime}$). For the other methods, we vary the number of clusters probed and the number of reranked documents. We perform a grid search over the query hyperparameters and report the results on the Pareto optimal hyperparameter combinations.
ColBERTv2 embeddings. First, we consider BEIR datasets embedded using the ColBERTv2 model (See Table 2, Fig. 4, and Appendix C.1). Lemur consistently outperforms the earlier methods. In particular, at 80% recall, it is 5-11x faster than the best-performing baseline method.
Other text embedding models. We then consider the BEIR datasets SCIDOCS (see Fig. 5) and Quora (see Appendix C.2) embedded using the answerai-colbert-small-v1, GTE-ModernColBERT-v1, LFM2-ColBERT-350M, and jina-colbert-v2 models. Again, Lemur significantly outperforms the baseline methods. We also observe that some of the baseline methods struggle on these embeddings. In particular, MUVERA fails to reach over 60% recall on the answerai-colbert-small, GTE-ModernColBERT, and LFM2-ColBERT embeddings.
Visual document models. Finally, we consider the ViDoRe dataset embedded using the ColModernVBERT and ColQwen2-v1.0 models (see Fig. 6). Lemur outperforms the baseline methods in both cases. Since our default training set selection strategy is not available for visual document retrieval models, we simply use documents embedded using the document encoder $D$ as a training set (the third strategy mentioned in Sec. 4.2); using a separate query set for training yields even better performance (see Appendix D.2).
Figure 5: End-to-end performance comparison using four different modern multi-vector text models on the SCIDOCS dataset. On all datasets, Lemur is significantly faster than the baseline methods, while especially MUVERA struggles on the non-ColBERTv2 models.

### v2

@@ 4 Implementation
In this section, we discuss how to implement the framework introduced in Section 3 in practice. First, we describe the network architecture we use. Then, we outline how model training can be scaled to large corpora by pretraining the feature encoder on a subset of the documents. Finally, we show that Lemur is robust to the choice of training data.
@@ 4.1 Network Architecture
We use a small two-layer MLP as the network $\phi$. This allows us to keep the inference cost trivial and to train the network quickly, even on a CPU. The feature encoder $\psi$ is defined as
$\psi(x)=\mathrm{LN}(\mathrm{GELU}(W^{\prime}x+b)),$
where $\mathrm{LN}(\cdot)$ is layer normalization (Ba et al., 2016) and $\mathrm{GELU}(\cdot)$ is the GELU activation function (Hendrycks and Gimpel, 2016), $W^{\prime}\in\mathbb{R}^{d^{\prime}\times d}$, and $b\in\mathbb{R}^{d^{\prime}}$.
We did not observe significant performance gains from increasing the network depth beyond two layers, since each $g_{l}$ is a convex piecewise-linear function. Based on our ablation study (see Section 6.2), we fix the hidden layer size to $d^{\prime}=2048$ in all other experiments. Increasing $d^{\prime}$ further improves the accuracy of the MaxSim estimates but has only a small impact on the end-to-end performance since it also increases the computational cost of retrieval.
In all experiments, we always use the same hyperparameters to train the model; see Appendix A.
4.2 Scalable Model Training
To scale model training to large corpora, we pretrain the feature encoder $\psi$ using a subset of the documents as targets. Specifically, we sample $m^{\prime}\ll m$ documents and define the target function $^{\prime}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m^{\prime}}, g=g$$^{\prime}=g_{I^{\prime}}$, where $I^{\prime}\subset[m]$ is the set of indices of the $m^{\prime}$ sampled documents. Using these auxiliary targets, we train a model $\phi^{\prime}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m^{\prime}}$,
$\phi^{\prime}(x)=W^{\prime\prime}\psi(x),$
where $W^{\prime\prime}\in\mathbb{R}^{m^{\prime}\times d^{\prime}}$.
We fix the weights of the feature encoder $\psi$ and learn the $j$th row of the output layer’s final weight matrix $W$ for the final model $\phi$ by solving
$w_{j}:=\underset{\beta\in\mathbb{R}^{d^{\prime}}}{\operatorname*{arg\,min}}\,\,\mathbb{E}\left[\left(\beta^{\top}\psi(x)-g_{j}(x)\right)^{2}\right]$ (7)
for each $j\in[m]$.
When the weights of $\psi$ are fixed, (7) is a standard linear regression task with a closed-form OLS solution. To speed up computation, we sample a smaller training set of size $n^{\prime}$ to compute the OLS solutions.
After computing the OLS solutions for all $m$ documents, the rows $w_{l}$ of the weight matrix $W$ can be stored in an ANNS index. Although we present CPU-only results in this paper, we note that both the indexing and query processes of Lemur are also easy to implement on a GPU, using, for example, CAGRA (Ootomo et al., 2024) for GPU-based ANNS. Algorithm 2 summarizes the indexing pipeline.
Indexing speed and index updates. In our experimental setup (see Section 6.1), on our largest dataset (8.8 million documents), training the feature encoder $\psi$ takes 20 minutes on a CPU. Computing the weight matrix $W$ of the final projection layer takes 70 minutes, and building the ANN graph takes 198 minutes, for a total of 4.8 hours. The time to compute $W$ scales linearly with the number of documents, while ANN index construction is often near $O(n\log n)$ for graph indexes. We note that we have not tuned our hyperparameters (see Appendix A) with respect to index construction time and that indexing can be sped up significantly with only a minor decrease in performance.
Compared to prior methods, Lemur is more scalable to large datasets because it does not require clustering all token embeddings (e.g., 600 million embeddings into $2^{18}$ clusters on our biggest dataset), a step that is infeasible without a GPU on large datasets. Moreover, Lemur indexes much lower-dimensional vectors than MUVERA (see Section 6.2). Since the feature encoder is kept fixed, new documents can be indexed quickly in Lemur by computing the document weights via linear regression and inserting them into the ANNS index (with graph-based ANNS methods such as HNSW supporting online updates).
Algorithm 2 Lemur Indexing Process
0: Corpus embedding sets $\{C_{j}\}_{j=1}^{m}$, training token embeddings $\{x_{i}\}_{i=1}^{n}$, number of auxiliary targets $m^{\prime}$, OLS sample size $n^{\prime}$
0: Feature encoder $\psi$, ANNS index over $\{w_{j}\}_{j=1}^{m}$
1: Sample $I^{\prime}\subset[m]$ with $|I^{\prime}|=m^{\prime}$
2: Define the auxiliary target function $g^{\prime}=g_{I^{\prime}}$
3: Train MLP $\phi^{\prime}(x)=W^{\prime\prime}\psi(x)$ to estimate $g^{\prime}(x)$
[…]
5:for all $j\in[m]$ do
6:  $y_{j}\leftarrow[g_{j}(x_{1}),\ldots,g_{j}(x_{n^{\prime}})]^{\top}\in\mathbb{R}^{n^{\prime}}$
7:  $w_{j}\leftarrow Z^{+}y_{j}$
8:end for
[…]
10:return $\psi$ and $\mathcal{J}$
@@ 4.3 Training Set Selection
Because only some of the datasets used in our experiments have a separate set of training queries, we use the corpus documents encoded with the query encoder to train Lemur in all experiments for consistency. Formally, we sample a small subset of $n^{\prime}\ll m$ corpus documents $\mathcal{C}_{1},\dots\mathcal{C}_{n^{\prime}}$. We encode them with the query encoder of the multi-vector model to obtain $X_{i}=Q(\mathcal{C}_{i})$ for each $i\in[n^{\prime}]$, and sample $n$ token embeddings for the training set from the set $\bigcup_{i=1}^{n^{\prime}}X_{i}$.
If there is a separate training set of queries $\mathcal{X}_{1},\dots,\mathcal{X}_{n^{\prime}}$ available, the query encoder of the multi-vector model can be used to generate the multi-vector representations $X_{1},\dots,X_{n^{\prime}}$, where $X_{i}=Q(\mathcal{X}_{i})$, and the training set can be sampled from $\bigcup_{i=1}^{n^{\prime}}X_{i}$. Our results (see Appendix H.2) indicate that using actual queries (when available) for training yields even higher performance.
Notably, even when the training set is sampled directly from the embeddings of the corpus $C_{1},\dots,C_{m}$ (encoded using the document encoder $D$), Lemur achieves consistent performance and still outperforms the baseline methods (see Appendix H.1). This training method requires neither additional data nor the use of an encoder. In summary, our results suggest that Lemur is robust with respect to the type of data used to train the model.
Random hidden layer weights. We hypothesize that the robustness of Lemur is because both query and document tokens are mapped by design by the late interaction model into a common semantic embedding space. As a result, although the training sets differ, they all provide samples from a similar underlying geometry.
In Appendix H.3, we perform an experiment in which we do not train the feature encoder $\psi$ at all; instead, we use an MLP with random weights in the hidden layer. Thus, the only component that adapts to the training data is the final projection layer. This is often referred to as an extreme learning machine (ELM) (Huang et al., 2006). Although performance degrades relative to training the feature encoder, Lemur still works well in this scenario, suggesting that its performance does not primarily depend on learning a representation highly specific to the training data. Instead, the hidden layer mainly serves to expand the input into a richer feature space. Different training sets may slightly change the distribution of examples, but do not change the underlying geometry to which the linear layer is fitted.
@@ 5.1 Multi-Vector Similarity Search
The retrieval system of ColBERT (Khattab and Zaharia, 2020) relies on token-level pruning of documents: For each query token, the most similar document tokens are first retrieved from $\bigcup_{j=1}^{m}C_{j}$, i.e., from the set of all document tokens. The documents containing the retrieved tokens are reranked by computing their MaxSim similarities to the query. PLAID (Santhanam et al., 2022a), DESSERT (Engels et al., 2023), EMVB (Nardini et al., 2024), and IGP (Bian et al., 2025) improve latency and memory efficiency of this basic scheme by adding different pruning and approximation steps. WARP (Scheerer et al., 2025) is an efficient retrieval engine for models trained specifically with the XTR (Lee et al., 2023) objective.
Like Lemur, MUVERA (Jayaram et al., 2024) reduces multi-vector similarity search to single-vector similarity search by generating single-vector encodings for queries and documents, after which any single-vector ANNS library can be used for retrieval. However, the representations of MUVERA are data-oblivious, whereas the representations of Lemur are learned in a supervised fashion by a model that is directly optimized for the task of interest. This makes Lemur representations more accurate at much lower dimensionality (see Section 6.2) and more robust across different embedding models (see Section 6.3).
5.2 Distilling MaxSim Scores
TCT-ColBERT (Lin et al., 2020, 2021) learns a single-vector dense retriever by distilling ColBERT’s MaxSim scores. BGE-M3 (Chen et al., 2024) also uses self-knowledge distillation between dense, sparse, and ColBERT-style multi-vector retrieval scores, but trains a unified embedding model rather than a corpus-specific search reduction. In contrast, LEMUR does not train a generic text encoder for queries and documents: it learns a corpus-specific search reduction for an existing multi-vector model. This makes Lemur a lightweight, task-agnostic search approximation that avoids expensive dense-retriever training and can be applied directly to existing late interaction models.
@@ 5.3 Reducing the Storage Cost of Multi-Vector Models
Compared to single-vector retrieval, multi-vector retrieval methods have both larger memory footprints and higher latencies because of the large number of embeddings that have to be stored and processed. There exist many recent works (e.g., Clavié et al., 2024; MacAvaney et al., 2025; Yan et al., 2025; Veneroso et al., 2025; Qin et al., 2026) that address the increased storage cost by decreasing the number of embeddings per document. These methods are orthogonal to our approach and can be combined with it.
@@ 6.1 Experimental Setup
Each experiment is run on a compute node with two 20-core Intel Xeon Gold 6230 (Cascade Lake) CPUs. Each CPU supports AVX-512 instructions, and we run each experiment on all 40 cores by parallelizing over queries.
@@
Figure 2: Ablation study on the effect of the hidden layer size $d^{\prime}$ on the performance of Lemur. Left: Comparison of $\text{Recall100@}k^{\prime}$ for three values of $d^{\prime}$ as a function of the candidate set size $k^{\prime}$. Right: Comparison of the end-to-end performance between different values of $d^{\prime}$ with ANNS and reranking included. While larger values of $d^{\prime}$ can yield more accurate estimates, the end-to-end performance gap is narrower due to increased ANNS complexity, yielding diminishing returns.
Datasets. For text models, we use eight retrieval datasets from the BEIR (Thakur et al., 2021) benchmark. Following Santhanam et al. (2022a) and Jayaram et al. (2024), we use the development set for MS MARCO and the test set for all other datasets. For visual document retrieval models, we pool together images and queries from the five English-language ViDoRe V3 (Loison et al., 2026) datasets. Table 1 lists the number of documents, the number of test queries, and the average number of embeddings per document for each dataset (using ColBERTv2 for the text datasets and ColQwen2 for ViDoRe). We use the value $k=100$ in all experiments in this section; for additional results with $k=10$, see Appendix F.2.
Embedding models. In our main experiments, we use the widely used ColBERTv2 (Santhanam et al., 2022b) model. In Section 6.3, we also measure end-to-end performance on the answerai-colbert-small-v1 (Clavié, 2025), GTE-ModernColBERT-v1 (Chaffin, 2025), LFM2-ColBERT-350M (Amini et al., 2025), and jina-colbert-v2 (Jha et al., 2024) text retrieval models, and on the visual document retrieval models ColModernVBERT (Teiletche et al., 2025) and ColQwen2-v1.0 (Faysse et al., 2025).
@@ 6.2 Ablation Study
[…]
We then study the effect of the latent-space dimension $d^{\prime}$ of the model on end-to-end latency (see Figure 2, right, and Appendix F). While increasing $d^{\prime}$ from 1024 to 2048 significantly reduces latency, there is a point of diminishing returns when $d^{\prime}$ is increased to 4096: even though a larger latent space dimension yields higher-quality candidates (as established in Figure 2, left), it also increases the computational cost of querying the ANNS index, which partially offsets the accuracy gains. This also explains why the performance gap to 10240-dimensional FDEs is even larger in the end-to-end setting. While using the value $d^{\prime}=4096$ leads to lower latency than $d^{\prime}=2048$ on some of the datasets (see Appendix F), the differences in performance are small. Hence, we use $d^{\prime}=2048$ in our end-to-end experiments to reduce memory consumption.
MaxSim correlation. In Appendix D, we measure Pearson’s correlation coefficient and Spearman’s rank correlation coefficient between Lemur’s estimates and the exact MaxSim similarities. Both correlation coefficients are high (>> 0.94) for each dataset.
6.3 End-to-End Performance
@@
In this section, we compare the end-to-end performance of Lemur against MUVERA (Jayaram et al., 2024), DESSERT (Engels et al., 2023), IGP (Bian et al., 2025), and PLAID (Santhanam et al., 2022a). We use recall to measure effectiveness, and queries per second (QPS) to measure the efficiency of the algorithms.
Lemur implementation. For Lemur, we use PyTorch for both training and inference for the model described in Section 4. As a single-vector similarity search library, we use Glass (Wang, 2025), an efficient implementation of HNSW (Malkov and Yashunin, 2018) with scalar quantization. We implement MaxSim reranking in C++.
Baseline implementations. For MUVERA, we use the official C++ implementation to generate the FDEs. To ensure a fair comparison, we also combine MUVERA’s FDEs with Glass and the same reranking implementation used for Lemur. For DESSERT and IGP, we use the official C++ implementations, and for PLAID, we use the Rust implementation in the fast-plaid library (Sourty, 2025).
Table 2: Best QPS at $\geq 80\%$ recall ($k=100$). A dash indicates that the method did not reach the specified recall.
@@
HotpotQA 426 22 37 10
NQ 869 79 107 38 16
Quora 4068 787 679 284 89
Touche-2020 1996 371 47 147 29
TREC-Covid 1753 57 34 109 23
FiQA-2018 2416 239 310 242 51
SCIDOCS 2591 391 320 285 85
Index hyperparameters. For both Lemur and the baseline methods, we fix the index hyperparameters to values recommended by the authors, except for DESSERT, for which we perform a grid search over $L\in\{32,64\},C\in\{5,7\}$. Specifically, for Lemur, we set the hidden layer size to $d^{\prime}=2048$; for MUVERA, as recommended by Jayaram et al. (2024), we set $R_{\text{reps}}=40$, $k_{\text{sim}}=6$, $d_{\text{proj}}=d$ and then apply a final projection to generate 10240-dimensional FDEs; for PLAID, DESSERT, and IGP, we set the number of clusters to $16\sqrt{n}$ (rounded down to the nearest power of two), where $n$ is the number of corpus token embeddings.
Query hyperparameters. For Lemur and MUVERA, we vary the candidate-list-size parameter (ef_search) of the ANNS library and the number of reranked documents ($k^{\prime}$). For the other methods, we vary the number of clusters probed and the number of reranked documents. We perform a grid search over the query hyperparameters and report results for the Pareto-optimal hyperparameter combinations.
Figure 4: End-to-end performance comparison using four modern multi-vector text models on the Quora dataset. Across all models, Lemur is significantly faster than the baseline methods, with MUVERA struggling especially on the non-ColBERTv2 models.
@@
[…]
In Appendix C, we use nDCG@$10$ on the BEIR datasets to verify that Lemur also preserves the retrieval quality of ColBERTv2 while maintaining low latency.
Other text embedding models. We then consider the BEIR datasets Quora (see Figure 4) and SCIDOCS (see Appendix G.2) embedded using the answerai-colbert-small-v1, GTE-ModernColBERT-v1, LFM2-ColBERT-350M, and jina-colbert-v2 models. Again, Lemur significantly outperforms the baseline methods. We also observe that some of the baseline methods struggle with these embeddings. In particular, MUVERA fails to exceed 60% recall on the answerai-colbert-small-v1, GTE-ModernColBERT, and LFM2-ColBERT embeddings.
Visual document models. Finally, we consider the ViDoRe dataset embedded using the ColModernVBERT and ColQwen2-v1.0 models (see Figure 5). Lemur outperforms the baseline methods in both cases. The performance gap is smaller than on the text datasets due to the proportionally much higher cost of the reranking stage caused by bigger document (image) representations. Since our default training set selection strategy is not available for visual document retrieval models, we simply use documents embedded using the document encoder $D$ as a training set (the third strategy mentioned in Section 4.3); using a separate query set for training yields even better performance (see Appendix H.2).
Long queries. While the default number of query vectors used in ColBERTv2 is 32, for long queries, e.g., in agentic applications, using more vectors can be preferable. To test the effectiveness of Lemur with longer queries, in Appendix E, we use the ArguAna dataset from BEIR with 300 vectors per query, following Santhanam et al. (2022b). Lemur still maintains both retrieval performance and MaxSim approximation quality.